The Bradley Corporation produces a product with the following costs as of July 1, 20X1: Material $5 per unit Labor 3 per unit Overhead 1 per unit Beginning inventory at these costs on July 1 was 4,200 units. From July 1 to December 1, 20X1, Bradley produced 14,400 units. These units had a material cost of $5, labor of $6, and overhead of $4 per unit. Bradley uses LIFO inventory accounting.

a. Assuming that Bradley sold 17,800 units during the last six months of the year at $20 each, what is its gross profit?

b. What is the value of ending inventory?
Answer #1

Answer a:

Beginning inventory on July 1:

Cost of inventory per unit = Material cost per unit + Labor cost per unit + Overhead cost per unit

= $5 + 3 + 1

= $9 per unit

From July 1 to December 1, 20X1:

Cost of inventory per unit = $5 +$6 + $4 = $15 per unit

Finished Goods Inventory- LIFO methood Beginning Inventory Cost per unit Total cost Unit Cost per unit Total cost UnitCost per unit Total cost Unit Production Cost of goods sold Ending Balance Date /Period Unit Cost per unit Total cost S9$37,800 July 1, 20X1 From July 1 to December 1, 20x1 Total 4,200 4,200.00 $9$37,800 $15 $216,000 14,400 3,400 $216,000 17,800 $15 $216,000 $9 $30,600 800 9 $7,200 14,400

As  Bradley Corporation uses LIFO:

Cost of goods sold as calculated above = 14,400 * $15 + 3,400 * $9 =$246,600

Sales value = Unit sold * sale price = 17,800 * $20 = $356,000

Gross profit = Sale - Cost of goods sold = $356,000 - $246,600 = $109,400

Gross profit = $109,400

Answer b:

Balance inventory from beginning inventory = 4,200 - 3,400 = 800 units

Value of ending inventory, as calculated in above table = 800 * $9 = $7,200

Value of ending inventory = $7,200